### v1

@@ Abstract
The all-pairs suffix-prefix (APSP) problem is a classical problem in string processing which has important applications in bioinformatics. Given a set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings, the APSP problem asks one to compute the longest suffix of $S_{i}$ that is a prefix of $S_{j}$ for all $k^{2}$ ordered pairs $\langle S_{i},S_{j}\rangle$ of strings in $\mathcal{S}$. In this paper, we consider the dynamic version of the APSP problem that allows for insertions of new strings to the set of strings. Our objective is, each time a new string $S_{i}$ arrives to the current set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings, to compute (1) the longest suffix of $S_{i}$ that is a prefix of $S_{j}$ and (2) the longest prefix of $S_{i}$ that is a suffix of $S_{j}$ for all $1\leq j\leq i$. We propose an $O(n)$-space data structure which computes (1) and (2) in $O(|S_{i}|\log\sigma+i)$ time for each new given string $S_{i}$, where $n$ is the total length of the strings.
Keywords:
all-pairs suffix-prefix problem Aho-Corasick automata directed acyclic word graphs (DAWGs)
@@ 1 Introduction
Before describing our APSP algorithm for the dynamic case (Section 5), we present a new AC-automata based APSP algorithm for a static set of strings (Section 4). While the algorithm of Loukides and Pissis [15] is based on a reduction of the APSP problem to the ULIT (union of labeled intervals on a tree) problem, our new APSP algorithm does not use this reduction. Instead, we use simple failure link traversal of the AC-automaton and marking operations on a compact prefix trie that is a compact version of the AC-trie. The simple nature of our algorithm makes it possible to work also on the DAWG structure in the dynamic case.
@@ 2.1 Strings
For a set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of strings, let $\mathsf{Substr}(\mathcal{S})=\bigcup_{i=1}^{k}\mathsf{Substr}(S_{i})$.
@@ 2.2 All-Pairs Suffix-Prefix Overlap (APSP) Problems
2.2.1 APSP Problems on Static Sets.
Problem 1 (static-APSP)
@@
Problem 2 (length-bounded static-APSP)
@@
2.2.2 APSP Problems on Dynamic Sets with Insertions.
Let $\mathcal{S}_{0}=\emptyset$ and for $1\leq i\leq k$, let $\mathcal{S}_{i}=\{S_{1},\ldots,S_{i}\}$ denote the subset of $\mathcal{S}$ consisting of the first $i$ strings. We consider a dynamic variant of APSP allowing for insertions of strings, named dynamic-APSP, that is defined as follows:
Problem 3 (dynamic-APSP)
The dynamic-APSP problem is, each time a new string $S_{i}$ is added to the current set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$, to compute the two following sets
@@
We remark that $\mathcal{L_{S}}=\bigcup_{i=1}^{k}\left(\mathcal{F}_{i}\cup\mathcal{B}_{i}\right)$ holds for $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$. Thus, a solution to Problem 3 is also a solution to Problem 1.
Problem 4 (length-bounded dynamic-APSP)
Let $\ell>0$ be a length threshold. The dynamic-APSP problem is, each time a new string $S_{i}$ is added to the current set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$, to compute the two following sets
@@
$\mathcal{B}_{i}^{\geq\ell}$ $=$ $\{\mathsf{lspo}(j,i)\mid 1\leq j\leq i,|\mathsf{lspo}(i,j)|\geq\ell\}.$
Note that a solution to Problem 4 is also a solution to Problem 2 as well.
@@ 3.3 Directed acyclic word graphs (DAWGs)
Figure 2: Illustration of $\mathsf{DAWG}(\mathcal{S})$ for the same set $\mathcal{S}=\{\mathrm{abaa,abac,abb,abcb,bab}$,
$\mathrm{babaa,bb,bbaa,bbba}\}$ of strings as in Fig. 1. The induced tree consisting only of the double-lined arcs is $\mathsf{Trie}(\mathcal{S})$.
@@ 4 Algorithm for Static-APSP
We build $\mathsf{AC}(\mathcal{S})$ by Theorem 3.1. We can compute $\mathsf{ComTrie}(\mathcal{S})$ and $\mathsf{des}(u)=v$ for all nodes $v$ in $\mathsf{Trie}(\mathcal{S})$ in time linear in the size of $\mathsf{Trie}(\mathcal{S})$, which is $O(n)$.
@@
We can reuse our data structure for different values of $\ell$, as we only need to change the underlined conditions in the above.
5 Algorithm for Dynamic-APSP
In this section, we present our algorithm for solving the dynamic-APSP problem: Given a new string $S_{i}$ that is inserted to the current set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings, our goal is to compute the outputs $\mathcal{F}_{i}$ and $\mathcal{B}_{i}$ of Problem 3.
@@ 5 Algorithm for Dynamic-APSP
Theorem 5.1
There exists a data structure of $O(n_{i})$ space, which solves the dynamic-APSP problem of computing $\mathcal{F}_{i}$ and $\mathcal{B}_{i}$ for a given new string $S_{i}$ in $O(|S_{i}|\log\sigma+i)$ time, where $n_{i}=\|\mathcal{S}_{i}\|=\sum_{j=1}^{i}|S_{i}|$.
Corollary 2
Let $\ell>0$ be a length threshold. There exists a data structure of $O(n_{i})$ space, which solves the length-bounded dynamic-APSP problem of computing $\mathcal{F}_{i}^{\geq\ell}$ and $\mathcal{B}_{i}^{\geq\ell}$ for a given new string $S_{i}$ in $O(|S_{i}|\log\sigma+i)$ time, where $n_{i}=\|\mathcal{S}_{i}\|=\sum_{j=1}^{i}|S_{i}|$.
We will focus on proving Theorem 5.1. Corollary 2 can be obtained from Theorem 5.1, in a similar manner to the static case in Section 4.
@@ 5.1 Computing $\mathcal{F}_{i}$
Now, we design our algorithm for the dynamic case with the DAWG structure, using the idea from Section 4 for the static case with the AC-automaton (see also Algorithm 1). Suppose that we have built and maintained the aforementioned data structures for the dynamic set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings. Given a new string $S_{i}$ to insert, we update $\mathsf{DAWG}(\mathcal{S}_{i-1})$ to $\mathsf{DAWG}(\mathcal{S}_{i})$ in $O(|S_{i}|\log\sigma)$ time with Theorem 3.2. We have now updated $\mathsf{SLTree}(\mathcal{S}_{i-1})$ to $\mathsf{SLTree}(\mathcal{S}_{i})$ as well.
@@
Lemma 1
There is a data structure of $O(n_{i})$ space, which computes $\mathcal{F}_{i}$ in $O(|S_{i}|\log\sigma+i)$ time per new string $S_{i}$ inserted.
Remark 1
@@ 5.2 Computing $\mathcal{B}_{i}$
Corollary 3
There is a data structure of $O(n_{i})$ space, which computes $\mathcal{B}_{i}$ in $O(|S_{i}|\log\sigma+i)$ time per new string $S_{i}$ inserted.

### v2

@@ Abstract
The all-pairs suffix-prefix (APSP) problem is a classical problem in string processing which has important applications in bioinformatics. Given a set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings, the APSP problem asks one to compute the longest suffix of $S_{i}$ that is a prefix of $S_{j}$ for all $k^{2}$ ordered pairs $\langle S_{i},S_{j}\rangle$ of strings in $\mathcal{S}$. In this paper, we consider the dynamic version of the APSP problem that allows for insertions of new strings to the set of strings. Our objective is, each time a new string $S_{i}$ arrives to the current set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings, to compute (1) the longest suffix of $S_{i}$ that is a prefix of $S_{j}$ and (2) the longest prefix of $S_{i}$ that is a suffix of $S_{j}$ for all $1\leq j\leq i$. We propose an $O(n)$-space data structure which computes (1) and (2) in $O(|S_{i}|\log\sigma+i)$ time for each new given string $S_{i}$, where $n$ is the total length of the strings and $\sigma$ is the alphabet size. Further, we show how to extend our methods to the fully dynamic version of the APSP problem allowing for both insertions and deletions of strings.
@@ 1 Introduction
Further, we consider the fully dynamic APSP problem which allows for both insertions and deletions of strings. We show how to solve fully dynamic APSP in $O(|S_{i}|\log\sigma+k)$ amortized time when string $S_{i}$ is inserted or deleted, where $k$ is the number of strings in the current set $\mathcal{S}$ of strings. In this fully dynamic version, we show how the DAWG-based algorithm can be simulated with the suffix tree in the case of insertions, and how our suffix-tree based method can handle the case of deletions. We emphasize that our approach is again quite different from Gusfield et al.’s method [10].
Before describing our APSP algorithms for the (fully) dynamic case (Section 5 and Section 6), we present a new AC-automata based APSP algorithm for a static set of strings (Section 4). While the algorithm of Loukides and Pissis [18] is based on a reduction of the APSP problem to the ULIT (union of labeled intervals on a tree) problem, our new APSP algorithm does not use this reduction. Instead, we use simple failure link traversals of the AC-automaton, and marking operations on a compact prefix trie that is a compact version of the AC-trie. The simple nature of our algorithm makes it possible to work also on the DAWG and on the suffix tree in the (fully) dynamic case.
@@
A preliminary version of this paper appeared in [13]. The new material in this full version is the algorithm for the fully dynamic APSP problem (Section 6), which was posed as an open question in the preliminary version [13].
@@ 2.1 Strings
For a set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of strings, let $\mathsf{Substr}(\mathcal{S})=\bigcup_{i=1}^{k}\mathsf{Substr}(S_{i})$. Let $\|\mathcal{S}\|=\sum_{i=1}^{k}|S_{i}|$ denote the total length of the strings in $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$.
@@ 2.2 All-Pairs Suffix-Prefix Overlap (APSP) Problems
2.2.1 APSP Problems on Static Sets
Problem 1 (static-APSP).
@@ 2.2.1 APSP Problems on Static Sets
Problem 2 (length-bounded static-APSP).
@@
2.2.2 APSP on Dynamic Sets with Insertions
We consider a dynamic variant of APSP allowing for insertions of strings, named dynamic APSP, that is defined as follows:
Problem 3 (dynamic APSP).
The dynamic APSP problem is, each time a new string $S_{i}$ is inserted to the current set $\mathcal{S}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings, to compute the two following sets
@@ 2.2.2 APSP on Dynamic Sets with Insertions
We remark that $\mathcal{L_{S}}=\bigcup_{i=1}^{k}\left(\mathcal{F}_{i}\cup\mathcal{B}_{i}\right)$ holds for the set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings. Thus, a solution to Problem 5 is also a solution to Problem 1.
Problem 4 (length-bounded dynamic APSP).
Let $\ell>0$ be a length threshold. The dynamic-APSP problem is, each time a new string $S_{i}$ is inserted to the current set $\mathcal{S}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings, to compute the two following sets
@@
$\mathcal{B}_{i}^{\geq\ell}$ $=$ $\{\mathsf{lspo}(j,i)\mid 1\leq j\leq i,|\mathsf{lspo}(j,i)|\geq\ell\}.$
Analogously, $\mathcal{L_{S}}^{\geq\ell}=\bigcup_{i=1}^{k}\left(\mathcal{F}_{i}^{\geq\ell}\cup\mathcal{B}_{i}^{\geq\ell}\right)$ holds for the set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of $k$ strings. Thus, a solution to Problem 6 is also a solution to Problem 2 as well.
2.2.3 APSP on Fully Dynamic Sets with Insertions and Deletions
We also consider a fully dynamic variant of APSP allowing for insertions and deletions of strings, named fully dynamic APSP, that is defined as follows:
Problem 5 (fully dynamic APSP).
In the fully dynamic APSP problem, when a new string $S_{i}$ is added to the current set $\mathcal{S}$ of strings, the goal is to compute the two following sets
$\mathcal{F}_{i}$ $=$ $\{\mathsf{lspo}(i,j)\mid j\in\mathsf{ind}(\mathcal{S})\},$
$\mathcal{B}_{i}$ $=$ $\{\mathsf{lspo}(j,i)\mid j\in\mathsf{ind}(\mathcal{S})\},$
and add them to the list $\mathcal{L}_{\mathcal{S}}$, where $\mathsf{ind}(\mathcal{S})$ denotes the set of indices $h$ such that $S_{h}\in\mathcal{S}$. When a string $S_{i}$ is deleted from the current set $\mathcal{S}$ of strings, then the goal is to remove $\mathcal{F}_{i}$ and $\mathcal{B}_{i}$ from the list $\mathcal{L}_{\mathcal{S}}$.
Problem 6 (length-bounded fully dynamic APSP).
Let $\ell>0$ be a length threshold. In the length-bounded fully dynamic APSP problem, when a new string $S_{i}$ is inserted to the current set $\mathcal{S}$ of strings, to compute the two following sets
$\mathcal{B}_{i}^{\geq\ell}$ $=$ $\{\mathsf{lspo}(j,i)\mid 1\leq j\leq i,|\mathsf{lspo}(j,i)|\geq\ell\},$
and add them to the list $\mathcal{L}_{\mathcal{S}}^{\geq\ell}$, where $\mathsf{ind}(\mathcal{S})$ denotes the set of indices $h$ such that $S_{h}\in\mathcal{S}$. When a string $S_{i}$ is deleted from the current set $\mathcal{S}$ of strings, then the goal is to remove $\mathcal{F}_{i}^{\geq\ell}$ and $\mathcal{B}_{i}^{\geq\ell}$ from the list $\mathcal{L}_{\mathcal{S}}^{\geq\ell}$.
@@ 3.3 Directed acyclic word graphs (DAWGs)
See Fig. 2 and 3 for a concrete example of the DAWG and its suffix links.
Figure 2: $\mathsf{DAWG}(\mathcal{S})$ for the same set $\mathcal{S}=\{\mathrm{abaa,abac,abb,abcb,bab,babaa,bb,bbaa,bbba}\}$ of strings as in Fig. 1. The induced tree consisting only of the double-lined arcs is $\mathsf{Trie}(\mathcal{S})$.
@@
3.4 Suffix trees
The suffix tree for a set $\mathcal{S}=\{S_{1},\ldots,S_{k}\}$ of strings, denoted $\mathsf{STree}(\mathcal{S})$, is a compacted trie such that
  • each edge is labeled by a non-empty substring in $\mathsf{Substr}(\mathcal{S})$;
  • the labels of the out-edges of each node begin with distinct characters;
  • every internal node is branching (i.e. has two or more children) or represents a suffix in $\mathsf{Suffix}(\mathcal{S})$;
  • every suffix in $\mathsf{Suffix}(\mathcal{S})$ is represented by a node.
For ease of explanations, we identify each node of $\mathsf{STree}(\mathcal{S})$ with the string that the node represents.
The suffix link of a node $v$ in $\mathsf{STree}(\mathcal{S})$ points to the node that represents the string $\mathsf{str}(v)[2..|\mathsf{str}(v)|]$.
See Fig. 4 and 5 for a concrete example of the suffix tree and its suffix links.
Figure 4: $\mathsf{STree}(\mathcal{S})$ for the same set $\mathcal{S}=\{\mathrm{abaa,abac,abb,abcb,bab,babaa,bb,bbaa,bbba}\}$ of strings as in Fig. 1. The induced tree consisting only of the double-lined arcs is a compacted version of $\mathsf{Trie}(\mathcal{S})$.
Figure 5: Illustration of the suffix links of $\mathsf{STree}(\mathcal{S})$ for the same set $\mathcal{S}$ of strings as in Fig 4.
Each edge label $x$ of $\mathsf{STree}(\mathcal{S})$ is represented by a tuple $(i,b,e)$ such that $x=S_{i}[b..e]$. This way $\mathsf{STree}(\mathcal{S})$ can be stored in $O(n)$ space, where $n=\|S\|$
Our version of suffix trees is the so-called Weiner tree [27] in which the repeated suffixes with unique right-extensions are represented by non-branching internal nodes (For example, string $\mathrm{bab}\in\mathsf{Suffix}(\mathcal{S})$ in Fig. 4 is represented by a non-branching internal node). This version of the suffix tree can easily be obtained by first building the suffix tree for strings $S_{1}\$_{1},\ldots,S_{k}\$_{k}$ with unique end-markers \$$\$_{1},\ldots,\$_{k}$ (\$$\$_{i}\neq\$_{j}$) and then by removing all edges labeled by the end-markers \$$\$_{1},\ldots,\$_{k}$.
By adapting Weiner’s construction [27] or Ukkonen’s construction [26] both working in online manners, one can update the suffix tree efficiently when a new string is inserted to the current set of strings. This setting is called as the semi-dynamic online construction of suffix trees in the literature [24].
Theorem 3 ([27, 26, 24]).
Suppose that $\mathsf{STree}(\mathcal{S}_{i-1})$ has been built for a set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings. Given a new string $S_{i}$ to insert, one can update $\mathsf{STree}(\mathcal{S}_{i-1})$ to $\mathsf{STree}(\mathcal{S}_{i})$ in $O(|S_{i}|\log\sigma)$ time.
@@ 4 Algorithm for Static-APSP
We build $\mathsf{AC}(\mathcal{S})$ by Theorem 1. We can compute $\mathsf{ComTrie}(\mathcal{S})$ and $\mathsf{des}(u)=v$ for all nodes $v$ in $\mathsf{Trie}(\mathcal{S})$ in time linear in the size of $\mathsf{Trie}(\mathcal{S})$, which is $O(n)$. ∎
@@
We can reuse our data structure for different values of $\ell$, as we only need to change the underlined conditions in the above. ∎
5 Algorithm for Dynamic APSP
In this section, we present our algorithm for solving the dynamic APSP problem: Given a new string $S_{i}$ that is inserted to the current set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings, our goal is to compute the outputs $\mathcal{F}_{i}$ and $\mathcal{B}_{i}$ of Problem 5.
@@ 5 Algorithm for Dynamic APSP
Theorem 5.
There exists a DAWG-based data structure of $O(n_{i})$ space, which solves the dynamic APSP problem of computing $\mathcal{F}_{i}$ and $\mathcal{B}_{i}$ for a given new string $S_{i}$ in $O(|S_{i}|\log\sigma+i)$ time, where $n_{i}=\|\mathcal{S}_{i}\|=\sum_{j=1}^{i}|S_{i}|$.
Corollary 2.
Let $\ell>0$ be a length threshold. There exists a DAWG-based data structure of $O(n_{i})$ space, which solves the length-bounded dynamic APSP problem of computing $\mathcal{F}_{i}^{\geq\ell}$ and $\mathcal{B}_{i}^{\geq\ell}$ for a given new string $S_{i}$ in $O(|S_{i}|\log\sigma+i)$ time, where $n_{i}=\|\mathcal{S}_{i}\|=\sum_{j=1}^{i}|S_{i}|$.
We will focus on proving Theorem 5. Corollary 2 can be obtained from Theorem 5, in a similar manner to the static case in Section 4.
@@ 5.1 Computing $\mathcal{F}_{i}$
Now, we design our algorithm for the dynamic case with the DAWG structure, using the idea from Section 4 for the static case with the AC-automaton (see also Algorithm 1). Suppose that we have built and maintained the aforementioned data structures for the dynamic set $\mathcal{S}_{i-1}=\{S_{1},\ldots,S_{i-1}\}$ of $i-1$ strings. Given a new string $S_{i}$ to insert, we update $\mathsf{DAWG}(\mathcal{S}_{i-1})$ to $\mathsf{DAWG}(\mathcal{S}_{i})$ in $O(|S_{i}|\log\sigma)$ time with Theorem 2. We have now updated $\mathsf{SLTree}(\mathcal{S}_{i-1})$ to $\mathsf{SLTree}(\mathcal{S}_{i})$ as well.
@@
Lemma 1.
There is a DAWG-based data structure of $O(n_{i})$ space, which computes $\mathcal{F}_{i}$ in $O(|S_{i}|\log\sigma+i)$ time per new string $S_{i}$ inserted.
Remark 1.
@@ 5.2 Computing $\mathcal{B}_{i}$
Corollary 3.
There is a DAWG-based data structure of $O(n_{i})$ space, which computes $\mathcal{B}_{i}$ in $O(|S_{i}|\log\sigma+i)$ time per new string $S_{i}$ inserted.
6 Algorithm for Fully Dynamic APSP
In this section, we present a suffix tree based algorithm for the fully dynamic APSP problem allowing for insertions and deletions of strings.
We make use of the following observation in our algorithm.
Observation 2.
The induced tree of $\mathsf{STree}(\mathcal{S})$ that consists only of the paths spelling out the strings in $\mathcal{S}$ is a compacted version of $\mathsf{Trie}(\mathcal{S})$. The set of nodes of $\mathsf{ComTrie}(\mathcal{S})$ is a subset of the nodes of this induced tree.
See Fig. 1 and Fig. 4 for a concrete example of this observation.
6.1 Insertions
In the case of insertions, one can use essentially the same algorithm from Section 5: When a new string $S_{i}$ is inserted to the current set $\mathcal{S}$, we first update the suffix tree in $O(|S_{i}|\log\sigma)$ time by Theorem 3. Then, we climb up the suffix link path from the node representing $S_{i}$ toward the root. By Observation 2, the $\mathsf{des}(\cdot)$ function to access the corresponding nodes in $\mathsf{ComTrie}(\mathcal{S})$ can be maintained analogously. In more detail, by Observation 2, a single edge in $\mathsf{ComTrie}(\mathcal{S})$ may be split into multiple edges in the compacted version of $\mathsf{Trie}(\mathcal{S})$. Thus, when an edge $(u,v)$ in the compacted version of $\mathsf{Trie}(\mathcal{S})$ is split into two edges $(u,w)$ and $(w,v)$ during the update of the suffix tree, then we simply set $\mathsf{des}(w)\leftarrow\mathsf{des}(v)$. These all give us how to compute $\mathcal{F}_{i}$ in $O(|S_{i}|\log\sigma)$ time using $\mathsf{STree}(\mathcal{S})$ of $O(n)$ space, where $n=\|\mathcal{S}\|$, as in Section 5.1.
One can compute $\mathcal{B}_{i}$ in $O(|S_{i}|\log\sigma)$ time and $O(n)$ space by using $\mathsf{STree}(\mathcal{S}^{R})$ for the set $\mathcal{S}^{R}$ of reversed strings, as in Section 5.2.
6.2 Deletions
We can also extend our suffix-tree based algorithm in the case where both insertions and deletions are supported.
Suppose that a string $S_{i}$ is deleted from the current set $\mathcal{S}$ of strings. Let $S_{i}[r..|S_{i}|]$ be the longest suffix of $S_{i}$ that corresponds to a branching node in the suffix tree, and let $S_{i}[t..|S_{i}|]$ be the longest suffix of $S_{i}$ such that $S_{i}[t..|S_{i}|]=S_{j}$ for some $j\in\mathsf{ind}(\mathcal{S})\setminus\{i\}$. Let $t=\infty$ when there is no such $j$ exists. Let $p=\min\{r,t\}$.
For each $1\leq q<p$ in increasing order, we remove the nodes $v_{q}$ (leaves or non-branching internal nodes) that respectively represent the suffixes $S_{i}[q..|S_{i}|]$. By tracing the suffix link chain from $v_{1}$ to $v_{p}$, this can be done in $O(1)$ time for each suffix $S_{i}[q..|S_{i}|]$ that is removed from the tree. If the parent of $v_{q}$ becomes non-branching after the removal of $v_{q}$, then the parent node is removed and its in-edge and out-edge are merged into a single edge. This is the same procedure as in the suffix tree in the sliding window model [15, 23, 16], and the total time complexity is $O(p\log\sigma)\subseteq O(|S_{i}|\log\sigma)$ for this step. Now the tree topology of the suffix tree has been updated.
Recall that each edge label $x$ is represented by a tuple $(i,b,e)$ such that $x=S_{i}[b..e]$. When $S_{i}$ is removed from $\mathcal{S}$, we need to replace a tuple $(i,b,e)$ which refers to $S_{i}$ with some tuple $(i^{\prime},b^{\prime},e^{\prime})$ such that $i^{\prime}\in\mathsf{ind}(\mathcal{S})\setminus\{i\}$ and $x=S_{i^{\prime}}[b^{\prime}..e^{\prime}]$. This can be done in $O(1)$ time by adapting the leaf-pointer algorithm [16] originally proposed for the sliding suffix trees. Leonard et al. [16] showed that edge-label maintenance in the suffix tree reduces to maintaining a pointer from each node to an arbitrary leaf in its subtree. They showed how to maintain such pointers in worst-case $O(1)$ time per leaf deletion. By applying this technique, we can update edge labels of $\mathsf{STree}(\mathcal{S})$ in a total of $O(p)\subseteq O(|S_{i}|)$ time.
Analogously, we update $\mathsf{STree}(\mathcal{S}^{R})$ in $O(|S_{i}|\log\sigma)$ time for the deleted string $S_{i}$.
In this fully dynamic setting, we maintain the output list $\mathcal{L}_{\mathcal{S}}$ with two arrays $F$ and $B$ such that $F[i]$ stores $\mathsf{lspo}(i,j)$ for all $j\in\mathsf{ind}(\mathcal{S})$ and $B[i]$ stores $\mathsf{lspo}(j,i)$ for all $j\in\mathsf{ind}(\mathcal{S})$. When a string $S_{i}$ is deleted, then we simply remove all elements in $F[i]$ and in $B[i]$. This clearly takes $O(k)$ time, where $k=|\mathcal{S}|$. We maintain these two dynamic arrays $F$ and $B$ by a standard doubling/halving technique: When all the entries of $F$ and $B$ become full, then we rebuild two arrays with $2|F|=2|B|$ entries each. When the number of occupied entries in $F$ and in $B$ reaches $|F|/4=|B|/4$, then we rebuild two arrays with $|F|/2=|B|/2$ entries each, and refresh the string ids. This takes $O(k)$ amortized time per insertion/deletion of a string, where $k$ is the number of strings in the set $\mathcal{S}$.
Finally, we obtain the following results:
Theorem 6.
There exists a suffix-tree-based data structure of $O(n)$ space, which solves the fully dynamic APSP problem in $O(|S_{i}|\log\sigma+k)$ amortized time for each inserted/deleted string $S_{i}$, where $n=\|\mathcal{S}\|$ and $k=|\mathcal{S}|$.
Corollary 4.
Let $\ell>0$ be a length threshold. There exists a suffix-tree-based data structure of $O(n)$ space, which solves the fully dynamic APSP problem in $O(|S_{i}|\log\sigma+k)$ amortized time for each inserted/deleted string $S_{i}$, where $n=\|\mathcal{S}\|$ and $k=|\mathcal{S}|$.